• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack (Techs)

    Log entry time 22:11:48 on November 05, 2009

    Entry number 297666

    keyword=Right Q2 status - OFF

    Yesterday, prior to closing the hall, the Right Q2 power supply dumped when it was at 1400 amps. After resetting the supply it dumped at about 50 amps. The supply was turned off for the night. We looked at it today. There seem to be a number of problems, none of them with obvious sources. The supply had a delay turning ON and OFF in remote mode. In local mode there was no delay. It seems the optocoupler in the remote control circuit was damaged. We replaced the optocouplers in the remote interface and the supply turns ON and OFF correctly. Some strange behaviors in the x-server/EDM/gateway/EPICS interfacing are adding to the array of troubles. There was some initial thought the SCR's were damaged, but this is not clear. It may be only a spurious feature in waveform of the current input to the supply. For now, the 480VAC to the supply has been locked out. We will continue working on it when the hall is available tomorrow.